A cow is tethered to a ring on outside wall of a barn 10 meters from the corners. The barn is 20 meters wide and 40 meters long. The rope is 50 meters in length. Find the cows grazing area. There is 2 answers to problem. Need to find both.
1
Expert's answer
2020-06-01T19:27:44-0400

First way:


A = "(\pi*50^2)/2"

B=C="(\pi*40^2)/4"

X=A+B*2= "(3.14*50^2)\/2 + ((3.14*40^2)\/4)*2"

X="3925+2512=6437m^2"


Second way:



A="(\pi*50^2)/2"

B="(\pi*40^2)/4"

C=D="(\pi*20^2)/4"

X=A+B+C*2="(3.14*50^2)/2+(3.14*40^2)/4+((\pi*20^2)/4)*2"

X=3925+1256+628=5809"m^2"